Fletcher Corporation is debating whether to convert its all-equity capital structure to one that is 40% debt. Currently, there are 2042 shares outstanding selling at $73 per share. EBIT is expected to remain at $14194 per year forever. The interest rate on new debt is 8%, and there are no taxes. Suppose that Fletcher goes through with the recapitalization, but you prefer the previous all-equity structure. You can unlever you position and re-create the original capital structure by selling a portion of your 100 shares and using that money to purchase the firm’s debt issue. How many of your 100 shares must you sell?
Answer #1

Current Capital Structure - 100% Equity

It is 100% equity, therefore EBIT is the earnings available to equity shareholders'.

Earning per share (EPS) = Earnings available to equity shareholders' / No. of shares

Or, EPS = $14,194 / 2042 shares = $6.95102840352 or $6.95

We hold 100 shares, therefore our current earnings are -

Earnings = EPS x No. of shares held = $6.95 x 100 = $695

Proposed Capital Structure - 60% Equity, 40% Debt

Current Value of Firm = 2042 shares x $73 per share = $149,066

To Change the capital structure the firm will borrow 40% of the total firm value and repurchase shares at the current market price.

Amount of shares to be bought = $149,066 x 40% = $59,626.40

No. of shares bought = $59,626.40 / $73 per share = 816 shares (Approx )

No. of shares remaining after repurchase = 2042 - 816 = 1,226 shares

New earnings available to Equity shareholders = EBIT - Interest = $14194 - $59,626.40 x 8% = $9,423.888

New EPS = $9,423.888 / 1226 = $7.6866949429 or $7.69

Unlever

Since, the firm will borrow 40% debt to lever the firm, we will sell 40% of our shares @current market price and purchase the debt issue for that amount to unlever our positiion.

Share sold = 100 x 40% = 40 shares

Amount received on sale = 40 x $73 = $2,920

We will purchase debt for $2,920 @ 8%.

Earnings on shares held = 60 x 7.69 = $461.40

Interest Earned = $2920 x 8% = $233.60

Total Earnings = $695

which is equal to the earnings before change in capital structure. So, to unlever you need to sell 40 shares.